In the laboratory, a general chemistry student measured the pH of a 0.507 M aqueous solution of acetic acid to be 2.504.  

Use the information she obtained to determine the Ka for this acid.  
Answer #1

Concentration of acetic acid = 0.507 M

pH = 2.504

[H+] = 10-pH

[H+] = 10-2.504

[H+] = 3.133 x 10-3 M

ICE table HAc (aq) \rightleftharpoons H+ (aq) Ac- (aq)
Initial conc. 0.507 M 0 0
Change -x +x +x
Equilibrium conc. 0.507 M - x +x +x

Ka = [H+]eq[Ac-]eq / [HAc]eq

Ka = [(x) * (x)] / (0.507 M - x)

Ka = (x2) / (0.507 M - x)

x = [H+] = 3.133 x 10-3 M

Ka = (3.133 x 10-3 M)2 / (0.507 M - 3.133 x 10-3 M)

Ka = 1.95 x 10-5 M